Using the elimination method how do i solve 6x-9y=25.5 7x-3x=-14
Answer by mananth(16946)

the second equation does not have two variables
7x-4x=-14
3x=-14
x= -14/3
Therefore
6x= -28
plug x in first equation
6x-9y=25.5
-28-9y=25.5
9y=-55.5
y= 55.5/9
use a calc to find the value